# Obfuscated Location Disclosure for Remote ID Enabled Drones

### v1

@@ 3.4 Requirements
  • R3: No Persistent Internet Connection on Drones. The method to obfuscate the location should work on the UAV independently from the availability of a persistent Internet connection onboard.
  • R4: Maximum Messages Generation Time. To keep compliance with the RID regulation, the time necessary to generate a RID message including an obfuscated location should not exceed $1$ second.
  • R5: Maximum Communication Overhead. The communication overhead resulting from the introduction of location obfuscation should not lead to message fragmentation, due to the low reliability of broadcast communications and limited energy budget. Thus, RID messages should not exceed the Maximum Transmission Unit (MTU) of WiFi, i.e., $2,304$ B [21].
@@ 4 Related Work
Several contributions in the current literature investigate location privacy issues and protection mechanisms in the context of vehicular networks. Some contributions provide location privacy via anonymity, as it can make tracking more difficult. One way to provide location privacy via anonymity is via pseudonyms [22][23], i.e., short-term identifiers that keep changing on a (regular) time basis. Note that if an adversary knows the location and the time before the pseudonym changes, they can link the old pseudonym to the new one, breaking location privacy.
@@
In the avionic domain, Sampigethaya et al. [24] propose to create areas dedicated to pseudonym swap where aircraft do not transmit messages, making it harder for a passive adversary to track aircraft based on wireless transmissions. This strategy cannot apply to our scenario, as drones should comply with RID. Svaigen et al. [25] propose creating dummy data and mobile entities. When applied to our context, these techniques would break compliance with RID and cause extra energy consumption.
@@
Location privacy can also be achieved via location obfuscation, i.e., perturbating the location disclosed to other parties to prevent precise location tracking [26]. Geo-Indistinguishability is a well-known location obfuscation method proposed by Andres et al.  [27], which protects the user’s location over multiple location disclosures. Such an approach also preserves location utility to LBSs. Geo-Indistinguishability was designed for low-frequency location disclosure with no time correlation. The achievable location privacy decreases when sending messages frequently [28], as passive attackers can reveal the actual location through inference attacks [9]. Thus, many contributions extend Geo-Indistinguishability to face such inference attacks in various scenarios. Chatzikokolakis et al. [8] focus on supporting location traces for Geo-Indistinguishability. Their proposal reduces the computation cost by reusing noise, providing enough privacy for an entire day. However, they only used a limited number (30) of daily queries. Hua et al. [29] focus on reporting obfuscated locations in rapid succession. The proposed mechanism can work in both a basic and advanced mode. Such proposals assume an external trusted service computes the obfuscated location, so it does not apply to our scenario. As discussed by Svaigen et al. in [22], another option includes not revealing the exact location but the one of an entity somewhere in an area. The security and utility of such a system depend on the size of the site. Another option discussed by Svaigen et al. in [22] is to use a protocol that works on a node-to-node basis, which is also not an option if the drone is not working in a swarm. Sampigethaya et al. [24] also propose using cooperative schemes, which is inapplicable to our problem. Other proposed mechanisms either require time and energy-consuming computations before using the system, such as the one by Zhang et al. [30], or have a chance to release the actual location, such as the one by Xiong et al. [31], which cannot fulfil our requirements. Xiao et al. [9, 32] specifically focus on the temporal correlation between location reports. They use a $\delta$-location set to protect the areas where the entity is most likely and develop an obfuscation method, namely Planar Isotropic Mechanism (PIM), which supports frequent location releases through a lightweight location obfuscation mechanism. Cao et al. [33] combine Geo-Indistinguishability with the proposal by Xiao et al. [9] and focus on obfuscating specific events, i.e. “moving between two specific places”. Through this strategy, an adversary cannot tell if an event has occurred. However, such a method cannot protect trajectory privacy. In the drone context, a recent work by Enayati et al. considered the application of Laplacian Geo-Indistinguishability to protect UAV location privacy in goods delivery applications and drone-IoT applications [34]. However, the authors do not consider compliance with RID regulations. To sum up, we cross-compare in Tab. I the described techniques and evaluate if and how they fulfil our requirements (Sec. 3.4).
@@
[8] 🌕 🌑 🌕 🌑
[9] 🌑 🌑 🌕 🌑
[11] 🌑 🌑 🌕 🌑
[22] 🌕 🌑 🌕 🌑
[24] 🌑 🌕 🌕 🌑
[27] 🌕 🌑 🌕 🌑
[29] 🌑 🌕 🌑 🌑
[30] 🌑 🌑 🌕 🌑
[31] 🌕 🌑 🌕 🌑
[33] 🌕 🌑 🌕 🌑
[34] 🌑 🌕 🌑
@@
Note that none of the available solutions fulfils all our requirements, at the same time, motivating our work. The proposals which fulfil most of our requirements are the ones by Xiao et al. [9][11][29], and [30]. Some approaches cannot be improved to deal with trajectory privacy (e.g., [11]), while others cannot run on drones due to heavy computational demands([29] and [30]). As for the proposal by [9] et al., it has not been conceived for 3D scenarios, it has not been tailored to constrained drones, and no considerations about their utility in drone-based scenarios have been provided. We integrate such an approach into our proposed solution as it achieves the lower bound of location privacy in equivalent 2-D scenarios. We will describe the integration process, extension, and evaluation in the following sections.
@@ 5 OLO-RID
5.1 Location Obfuscation on RID-enabled Drones
Our solution, namely OLO-RID, leverages the mathematical model of the PIM approach proposed by Xiao et al. in [9]. We first review the original approach and then discuss its limitations and our contributions.
Introduction to PIM. For ease of discussion, consider a two-coordinate system (latitude-longitude). We divide the area around the moving entity into multiple cells $s_{1},s_{2},...,s_{N}\in S$. Based on previous travel data of the mobile entity or public data about such a journey, PIM requires us to create a transition matrix $\mathbf{M}$, holding the probability that the mobile entity moves from one cell to another. We denote $M_{ij}$ as the probability that the mobile entity moves from cell $i$ to cell $j$. We also consider a probability array $\mathbf{p}$, indicating the probability that the moving entity is in the cell $s_{i}$. As an example, if the mobile entity has a uniform probability of residing in cells 1, 2, 3, and 4, then $p=[0.25,0.25,0.25,0.25,0,0,...]$. We define $p_{t}^{-}$ as the prior probability at timestamp $t$, i.e., the probability that at time $t$, before a location release, the moving entity resides in a given cell. We also define $p_{t}^{+}$ as the posterior probability, i.e., the probability that at time $t$, after a location release, the moving entity resides in a given cell. According to PIM, we compute the prior probability as $p_{t}^{-}=p_{t-1}^{+}\mathbf{M}$. We can also compute the posterior probability according to Eq. 1, where $\mathbf{z}_{t}$ is the location released at time $t$, and $\mathbf{u}_{t}$ is the actual location.
@@ 5.1 Location Obfuscation on RID-enabled Drones
Consider the actual location of the mobile entity $\mathbf{x}$. The PIM mechanism aims to generate the obfuscated location offset $\mathbf{z^{\#}}$, s.t. $\mathbf{z}=\mathbf{x}+\mathbf{z^{\#}}$. To this aim, PIM applies the following procedure. We compute the convex hull from the $\delta$-location set $\Delta X$. Given a set of points and drawing lines between each point, the convex hull is the smallest area not crossing any line [9]. Denote a vertex of the convex hull as $v_{i}$. We denote $\Delta V$ as the set of all values $v_{i}-v_{j}$ for all possible $i$ and $j$. From this, we calculate the sensitivity hull $K$ of a convex hull of $\Delta V$. A sensitivity hull is a convex hull capturing the sensitivity between two releases from two instances, i.e., the difference between two queries if one piece of data has changed. From the sensitivity hull $K$, we sample points uniformly. These points are used in an isotropic transformation, resulting in the matrix $T$. To calculate the isotropic position from $K$, we get $\mathbf{K_{I}}=TK$. We sample from $\mathbf{K_{I}}$ a point uniformly at random, namely $\mathbf{z^{{}^{\prime}}}$. To increase randomness, we use a Gamma distribution $\Gamma(3,\epsilon^{-1})$ to get $r$. We denote the new random location as $\mathbf{z^{*}}=r\cdot\mathbf{z^{{}^{\prime}}}$. This point has to be transformed back to the original space $\mathbf{z^{\#}}=T^{-1}\mathbf{z^{*}}$. Finally, we can calculate our obfuscated location $\mathbf{z}$ using the real location $\mathbf{x}$ and $\mathbf{z^{\#}}$, thus we get $\mathbf{z}=\mathbf{x}+\mathbf{z^{\#}}$. Expanding the formula, we have that $\mathbf{z}=\mathbf{x}+rT^{-1}\mathbf{z^{{}^{\prime}}}$.
@@
Modifications to PIM. As discussed in Sec. 4, the solution introduced by Xiao et al. in [9], namely PIM, does not fulfill all our requirements natively, preventing its straightforward integration for our problem. First, PIM protects the location in a two-dimensional coordinate system. However, UAVs move across a three-dimensional space. RID messages still leak information when obfuscating only two dimensions, e.g., if the UAV is taking off or landing, leading to privacy issues. Thus, we extend the method by Xiao et al. to support location obfuscation in a 3-D space. Another downside of PIM when adopted in our context is the transition matrix, which contains a specific bounded area where the UAV is supposed to fly. When the UAV flies far outside these bounds, the computed obfuscated location is still based on a location within bounds, resulting in the computation of an obfuscated location very far from the actual location. In turn, such difference makes disclosed locations largely unrealistic. A naive solution could be uploading an extensive transition matrix on the UAV, so it likely always stays within bounds. However, a UAV is a constrained device with limited memory and computation power. Thus, loading an extensive transition matrix in memory is not an option. The computation time would also increase because of more extensive matrix computations. To solve this issue, our solution uses a small transition matrix supporting the areas around the UAV in all three dimensions. The drone gets information about the prior probabilities from publicly accessible or personal flight data. This is indeed feasible, as drones likely operate in well-defined areas, with limited size. In some cases, e.g., autonomous drones, the precise travel path is even known in advance. To generate the probabilities of a UAV flying from one cell to another, we initialise all the cells in the transition matrix to have the same probability $\frac{1}{b^{3}}$, in line with the logic of remotely piloted systems. Specifically, we check whether a cell near the UAV at timestamp $t$ is also present at timestamp $t-1$. If true, the cell is regularly updated as $p_{t}^{-}=M\cdot p_{t-1}^{+}$. Otherwise, the value of $p_{t}^{-}=M\cdot\frac{1}{b^{3}}$. Algo. 1 provides the pseudo-code of our modified location obfuscation technique.
@@ 15       Derive posterior probability $p_{t}^{+}$ by Equation 1;
We remark that the probability that the drone re-initializes the values of the prior probabilities to the default value ($\frac{1}{b^{3}}$) can be made arbitrarily small based on the knowledge of the area where the UAV has to fly and the autonomy of the battery deployed onboard.
@@ 5.2 OLO-RID Protocol flow
OLO-RID involves a Registration Phase, executed before drone deployment, and a Runtime Phase, executed during deployment. Fig. 1 provides an overview of the operations required in such two phases, as described below.
@@
    The Encryption Module encrypts the location $\mathbf{x}$ using the ECIES scheme (Sec. 2) as $c=PKE.ENC(x_{lat}||x_{lon}||x_{alt},pk_{USS})$, and sends it back to the RID Module.
@@
    The RID Module generates a RID message $m$ using the UID, obfuscated location $\mathbf{z}$, velocity $\mathbf{v}$, CS’s location $\mathbf{x_{CS}}$, the current timestamp $t$, the emergency status at time $t$ $em_{t}$, and the encrypted location $c$, as $m=UID||\mathbf{z}||\mathbf{v}||x_{CS}||t||em_{t}||c$.
@@ 5.3 Application of OLO-RID in Reference Use Cases
In this section, we describe how the services connected to the use cases described in Sec. 3.3 can use the location disclosed by the drones via OLO-RID.
@@ 6 Security Analysis
Location Obfuscation Privacy. Our solution integrates the approach proposed by Xiao et al. in [9] to disclose an obfuscated location $\mathbf{z}$ in place of the current UAV location $\mathbf{x}$. As noted by relevant works on location privacy  [8], achieving complete location privacy when frequently disclosing the current location is not feasible. However, as demonstrated by Xiao et al. in [9], the proposed solution achieves the lower bound of trajectory privacy when considering multiple time-correlated location disclosures.
As per the introduced modifications to the method by Xiao et al., we first notice that we apply location obfuscation on a 3-D space. As we use the algorithm independently on the horizontal (latitude and longitude) and vertical coordinates (altitude), such a change does not impact the location privacy of the single coordinates. Our proposed OLO-RID solution also modifies the computation of the prior probability array ($p_{t}^{-})$. Specifically, we generate all permutations of $[0,1,-1]$ and create a box around the current UAV location, consisting of $b^{3}$ cells, where $b$ denotes the number of cells in each dimension. Recall that UAVs can move in space across consecutive location disclosures. Thus, such a movement may lead to the generation of new cells if the UAV moves into a new cell. In OLO-RID, we minimize the probability of this event through the knowledge of the area where the drone is flying. When this change occurs, OLO-RID re-initialize all cells to $p_{\#}=\frac{1}{b^{3}}$. As per [9], we compute the prior probability as $p_{t}^{-}=p_{\#}M$, being $M$ the transition matrix. We recall the definition of adversarial privacy [9].
@@
In our setting, adversarial privacy is equivalent to differential privacy on the $\delta$-location, as shown by Xiao et al. [9]. Thus, we can re-use the same proof in Xiao et al. [9].
@@
It is equivalent to proving adversarial privacy on a 0-location set, which includes all possible locations. If $x_{t}^{*}\in\Delta X_{t}$, then $z_{t}$ is generated by $x_{t}^{*}$. As shown by Xiao et al. [9], $z_{t}$ is $\epsilon_{t}$-differentially private. So $\frac{Pr(u_{t}^{*}=s_{i}|z_{t})}{Pr(u_{t}^{*}=s_{t})}\leq e^{\epsilon}$. When $x_{t}^{*}\notin\Delta X_{t}$, we can replace $x_{t}^{*}$ with a surrogate $\tilde{x}_{t}$, and thus:
@@
Encrypted Location Report Robustness. OLO-RID requires delivering encrypted location reports, i.e., cryptography values allowing the TTP (USS) to obtain the plain-text drone location, if necessary. To do so, we use the ECIES public-key encryption scheme, implying the generation of an ephemeral symmetric key signed through the public key of the USS. Such location reports can only be decrypted using the private key of the USS, which is assumed to stay private on the USS throughout the deployment of the system. If the private key of the USS is not leaked, based on the cryptographic properties of ECIES, a passive eavesdropper cannot obtain the plain-text location $\mathbf{x}$ used to generate the ciphertext $c$. The eavesdropper cannot even know if the plain-text location $\mathbf{x}$ changes or stays the same across consecutive location disclosures, thanks to the protection offered by ECIES against chosen-ciphertext, and chosen-plaintext attacks, as discussed in [15].

### v2

@@ 3.4 Requirements
  • R3: No Persistent Internet Connection on UAVs. The method to obfuscate the location must work on the UAV independently from the availability of a persistent Internet connection onboard.
  • R4: Maximum Messages Generation Time. To keep compliance with the RID regulation, the time necessary to generate a RID message including an obfuscated location must not exceed $1$ second.
  • R5: Maximum Communication Overhead. The communication overhead resulting from the introduction of location obfuscation should not lead to message fragmentation, due to the low reliability of broadcast communications and limited energy budget. Thus, RID messages must not exceed the Maximum Transmission Unit (MTU) of WiFi, i.e., $2,304$ B [ieee2012].
@@ 4 Related Work
Several contributions in the current literature investigate location privacy issues and protection mechanisms in the context of vehicular networks. Some contributions provide location privacy via anonymity, as it can make tracking more difficult. One way to provide location privacy via anonymity is via pseudonyms [svaigen2022_iotmag][tedeschi2021_acsac], i.e., short-term identifiers that keep changing on a (regular) time basis. Note that if an adversary knows the location and the time before the pseudonym changes, they can link the old pseudonym to the new one, breaking location privacy.
@@
In the avionic domain, Sampigethaya et al. [Sampigethaya2009_dasc] propose to create areas dedicated to pseudonym swap where aircraft do not transmit messages, making it harder for a passive adversary to track aircraft based on wireless transmissions. This strategy cannot apply to our scenario, as UAVs should comply with RID. Svaigen et al. [svaigen2022_icc] propose creating dummy data and mobile entities. When applied to our context, these techniques would break compliance with RID and cause extra energy consumption.
@@
Location privacy can also be achieved via location obfuscation, i.e., perturbating the location disclosed to other parties to prevent precise location tracking [jiang2021_csur]. Geo-Indistinguishability is a well-known location obfuscation method proposed by Andres et al.  [andres2013_ccs], which protects the user’s location over multiple location disclosures. Such an approach also preserves location utility to LBSs. Geo-Indistinguishability was designed for low-frequency location disclosure with no time correlation. The achievable location privacy decreases when sending messages frequently [mendes2020_pets], as passive attackers can reveal the actual location through inference attacks [xiao2015_ccs]. Thus, many contributions extend Geo-Indistinguishability to face such inference attacks in various scenarios. Chatzikokolakis et al. [chatzikokolakis2014_pets] focus on supporting location traces for Geo-Indistinguishability. Their proposal reduces the computation cost by reusing noise, providing enough privacy for an entire day. However, they only used a limited number (30) of daily queries. Hua et al. [hua2017_tifs] focus on reporting obfuscated locations in rapid succession. The proposed mechanism can work in both a basic and advanced mode. Such proposal assumes an external trusted service computes the obfuscated location, so it does not apply to our scenario. As discussed by Svaigen et al. in [svaigen2022_iotmag], another option includes not revealing the exact location but the one of an entity somewhere in an area. The security and utility of such a system depend on the size of the site. Another option discussed by Svaigen et al. in [svaigen2022_iotmag] is to use a protocol that works on a node-to-node basis, which is also not an option if the UAV is not working in a swarm. Sampigethaya et al. [Sampigethaya2009_dasc] also propose using cooperative schemes, which is inapplicable to our problem. Other proposed mechanisms either require time and energy-consuming computations before using the system, such as the one by Zhang et al. [zhang2019_tifs], or have a chance to release the actual location, such as the one by Xiong et al. [xiong2019_ijdsn], which cannot fulfil our requirements. Xiao et al. [xiao2015_ccs, xiao2017_vldb] specifically focus on the temporal correlation between location reports. They use a $\delta$-location set to protect the areas where the entity is most likely and develop an obfuscation method, namely Planar Isotropic Mechanism (PIM), which supports frequent location releases through a lightweight location obfuscation mechanism. Cao et al. [cao2019_icde] combine Geo-Indistinguishability with the proposal by Xiao et al. [xiao2015_ccs] and focus on obfuscating specific events, i.e. “moving between two specific places”. Through this strategy, an adversary cannot tell if an event has occurred. However, such a method cannot protect trajectory privacy. In the UAV context, a recent work by Enayati et al. considered the application of Laplacian Geo-Indistinguishability to protect UAV location privacy in goods delivery applications and UAV-IoT applications [enayati2023_iotj]. However, the authors do not consider compliance with RID regulations. To sum up, we cross-compare in Tab. I the described techniques and evaluate if and how they fulfil our requirements (Sec. 3.4).
@@
[chatzikokolakis2014_pets] 🌕 🌑 🌕 🌑
[xiao2015_ccs] 🌑 🌑 🌕 🌑
[brighente2022_ares] 🌑 🌑 🌕 🌑
[svaigen2022_iotmag] 🌕 🌑 🌕 🌑
[Sampigethaya2009_dasc] 🌑 🌕 🌕 🌑
[andres2013_ccs] 🌕 🌑 🌕 🌑
[hua2017_tifs] 🌑 🌕 🌑 🌑
[zhang2019_tifs] 🌑 🌑 🌕 🌑
[xiong2019_ijdsn] 🌕 🌑 🌕 🌑
[cao2019_icde] 🌕 🌑 🌕 🌑
[enayati2023_iotj] 🌑 🌕 🌑
@@
Note that none of the available solutions fulfils all our requirements, at the same time, motivating our work. The proposals which fulfil most of our requirements are the ones by Xiao et al. [xiao2015_ccs][brighente2022_ares][hua2017_tifs], and [zhang2019_tifs]. Some approaches cannot be improved to deal with trajectory privacy (e.g., [brighente2022_ares]), while others cannot run on UAVs due to heavy computational demands ([hua2017_tifs] and [zhang2019_tifs]). As for the proposal by [xiao2015_ccs] et al., it has not been conceived for 3D scenarios, it has not been tailored to constrained UAVs, and no considerations about their utility in UAV-based scenarios have been provided. We integrate such an approach into our proposed solution as it achieves the lower bound of location privacy in equivalent 2-D scenarios. We will describe the integration process, extension, and evaluation in the following sections.
@@ 5 OLO-RID
This section provides the details of OLO-RID. We first describe the location obfuscation procedure in Sec. 5.1. Then, we describe the operations occurring on the UAV in Sec. 5.2 and, finally, provide the details of the additional procedures required for each of the use cases in Sec. 5.3.
5.1 Location Obfuscation on RID-enabled UAVs
Our solution, namely OLO-RID, leverages the mathematical model of the PIM approach proposed by Xiao et al. in [xiao2015_ccs]. We first review the original approach and then discuss its limitations and our contributions.
Introduction to PIM. For ease of discussion, consider a two-coordinate system (latitude-longitude). We divide the area around the moving entity into multiple cells $s_{1},s_{2},...,s_{N}\in S$. Based on previous travel data of the mobile entity or public data about such a journey, PIM requires us to create a transition matrix $\mathbf{M}$, holding the probability that the mobile entity moves from one cell to another. We denote $M_{ij}$ as the probability that the mobile entity moves from cell $i$ to cell $j$. We also consider a probability array $\mathbf{p}$, indicating the probability that the moving entity is in the cell $s_{i}$. As an example, if the mobile entity has a uniform probability of residing in cells 1, 2, 3, and 4, then p=[0.25,0.25,0.25,0.25,$0,$]p=[0.25,0.25,0.25,0.25,$0,$...]. We define $p_{t}^{-}$ as the prior probability at timestamp $t$, i.e., the probability that at time $t$, before a location release, the moving entity resides in a given cell. We also define $p_{t}^{+}$ as the posterior probability, i.e., the probability that at time $t$, after a location release, the moving entity resides in a given cell. According to PIM, we compute the prior probability as $p_{t}^{-}=p_{t-1}^{+}\mathbf{M}$. We can also compute the posterior probability according to Eq. 1, where $\mathbf{z}_{t}$ is the location released at time $t$, and $\mathbf{u}_{t}$ is the actual location.
@@ 5.1 Location Obfuscation on RID-enabled UAVs
Consider the actual location of the mobile entity $\mathbf{x}$. The PIM mechanism aims to generate the obfuscated location offset $\mathbf{z^{\#}}$, s.t. $\mathbf{z}=\mathbf{x}+\mathbf{z^{\#}}$. To this aim, PIM applies the following procedure. We compute the convex hull from the $\delta$-location set $\Delta X$. Given a set of points and drawing lines between each point, the convex hull is the smallest area not crossing any line [xiao2015_ccs]. Denote a vertex of the convex hull as $v_{i}$. We denote $\Delta V$ as the set of all values $v_{i}-v_{j}$ for all possible $i$ and $j$. From this, we calculate the sensitivity hull $K$ of a convex hull of $\Delta V$. A sensitivity hull is a convex hull capturing the sensitivity between two releases from two instances, i.e., the difference between two queries if one piece of data has changed. From the sensitivity hull $K$, we sample points uniformly. These points are used in an isotropic transformation, resulting in the matrix $T$. To calculate the isotropic position from $K$, we get $\mathbf{K_{I}}=TK$. We sample from $\mathbf{K_{I}}$ a point uniformly at random, namely $\mathbf{z^{{}^{\prime}}}$. To increase randomness, we use a Gamma distribution $\Gamma(3,\epsilon^{-1})$ to get $r$. We denote the new random location as $\mathbf{z^{*}}=r\cdot\mathbf{z^{{}^{\prime}}}$. Note that there is a small probability that the location $\mathbf{z^{*}}$ lies outside of the $\delta$-location set. In such cases, PIM uses a surrogate location $\tilde{x}_{t}$, selected as the cell in $\Delta X$ closer to the actual location $\mathbf{x}$. This location has to be transformed back to the original space $\mathbf{z^{\#}}=T^{-1}\mathbf{z^{*}}$. Finally, we can calculate our obfuscated location $\mathbf{z}$ using the real location $\mathbf{x}$ and $\mathbf{z^{\#}}$, thus we get $\mathbf{z}=\mathbf{x}+\mathbf{z^{\#}}$. Expanding the formula, we have that $\mathbf{z}=\mathbf{x}+rT^{-1}\mathbf{z^{{}^{\prime}}}$.
@@
Extension of PIM. As discussed in Sec. 4, the solution introduced by Xiao et al. in [xiao2015_ccs], namely PIM, does not fulfill all our requirements natively, preventing its straightforward integration for our problem. First, PIM protects the location in a two-dimensional coordinate system. However, UAVs move across a three-dimensional space. RID messages still leak information when obfuscating only two dimensions, e.g., if the UAV is taking off or landing, leading to privacy issues. Thus, we extend the method by Xiao et al. to support location obfuscation in a 3-D space. Another downside of PIM when adopted in our context is the transition matrix, which contains a specific bounded area where the UAV is supposed to fly. When the UAV flies far outside these bounds, the computed obfuscated location is still based on a location within bounds, resulting in the computation of an obfuscated location very far from the actual location. In turn, such difference makes disclosed locations largely unrealistic. A naive solution could be uploading an extensive transition matrix on the UAV, so it likely always stays within bounds. However, a UAV is a constrained device with limited memory and computation power. Thus, loading an extensive transition matrix in memory is not an option. The computation time would also increase because of more extensive matrix computations. To solve this issue, our solution uses a small transition matrix supporting the areas around the UAV in all three dimensions. The UAV gets information about the prior probabilities from publicly accessible or personal flight data. This is indeed feasible, as UAVs likely operate in well-defined areas, with limited size. In some cases, e.g., autonomous UAVs, the precise travel path is even known in advance. However, in the more general case of RPASs, a UAV could move in any direction with the same probability. To generate the probabilities of a UAV flying from one cell to another, we initialise all the cells in the transition matrix to have the same probability $\frac{1}{b^{3}}$, in line with the logic of remotely piloted systems. Specifically, we check whether a cell near the UAV at timestamp $t$ is also present at timestamp t1t-1. If true, the cell is regularly updated as $p_{t}^{-}=p_{t-1}^{+}\cdot M$. Otherwise, the value of $p_{t}^{-}=\frac{1}{b^{3}}\cdot M.$ Algo. 1 provides the pseudo-code of our modified location obfuscation technique.
@@ 15   Derive posterior probability $p_{t}^{+}$ by Equation 1;
We remark that the probability that the UAV re-initializes the values of the prior probabilities to the default value ($\frac{1}{b^{3}}$) can be made arbitrarily small based on the knowledge of the area where the UAV has to fly and the autonomy of the battery deployed onboard. Finally, we integrate the extended mechanism discussed above into the RID specification, defining the complete message format and protocol flow in a way that is compliant and applicable to RID-enabled devices, as well as backward compatible with RID receivers.
@@ 5.2 OLO-RID Protocol flow
OLO-RID involves a Registration Phase, executed before UAV deployment, and a Runtime Phase, executed during deployment. Fig. 1 provides an overview of the operations required in such two phases, as described below.
@@
[…]
@@
[…]
@@ 5.3 Application of OLO-RID in Reference Use Cases
In this section, we describe how the services connected to the use cases described in Sec. 3.3 can use the location disclosed by the UAVs via OLO-RID.
@@ 6 Security Analysis
Location Obfuscation Privacy. Our solution integrates the approach proposed by Xiao et al. in [xiao2015_ccs] to disclose an obfuscated location $\mathbf{z}$ in place of the current UAV location $\mathbf{x}$. As noted by relevant works on location privacy  [chatzikokolakis2014_pets], achieving complete location privacy when frequently disclosing the current location is not feasible. However, as demonstrated by Xiao et al. in [xiao2015_ccs], the proposed solution achieves the lower bound of trajectory privacy when considering multiple time-correlated location disclosures.
As per the introduced modifications to the method by Xiao et al., we first notice that we apply location obfuscation on a 3-D space. As we use the algorithm independently on the horizontal (latitude and longitude) and vertical coordinates (altitude), such a change does not impact the location privacy of the single coordinates. Our proposed OLO-RID solution also modifies the computation of the prior probability array ($p_{t}^{-})$. Specifically, we generate all permutations of [0,1,1][0,1,-1] and create a box around the current UAV location, consisting of $b^{3}$ cells, where $b$ denotes the number of cells in each dimension. Recall that UAVs can move in space across consecutive location disclosures. Thus, such a movement may lead to the generation of new cells if the UAV moves into a new cell. In OLO-RID, we minimize the probability of this event through the knowledge of the area where the UAV is flying. When this change occurs, OLO-RID re-initialize all cells to $p_{\#}=\frac{1}{b^{3}}$. As per [xiao2015_ccs], we compute the prior probability as $p_{t}^{-}=p_{\#}M$, being $M$ the transition matrix. We recall the definition of adversarial privacy [xiao2015_ccs].
@@
In our setting, adversarial privacy is equivalent to differential privacy on the $\delta$-location, as shown by Xiao et al. [xiao2015_ccs]. Thus, we can apply the same reasoning provided by Xiao et al. in [xiao2015_ccs] to prove the security of our solution, as follows.
@@
It is equivalent to proving adversarial privacy on a 0-location set, which includes all possible locations. If $x_{t}^{*}\in\Delta X_{t}$, then $z_{t}$ is generated by $x_{t}^{*}$. As shown by Xiao et al. [xiao2015_ccs], $z_{t}$ is $\epsilon_{t}$-differentially private. So $\frac{Pr(u_{t}^{*}=s_{i}|z_{t})}{Pr(u_{t}^{*}=s_{t})}\leq e^{\epsilon}$. When $x_{t}^{*}\notin\Delta X_{t}$, we can replace $x_{t}^{*}$ with a surrogate $\tilde{x}_{t}$, and thus:
@@
Encrypted Location Report Robustness. OLO-RID requires delivering encrypted location reports, i.e., cryptography values allowing the TTP (USS) to obtain the plain-text UAV location, if necessary. To do so, we use the ECIES public-key encryption scheme, implying the generation of an ephemeral symmetric key signed through the public key of the USS. Such location reports can only be decrypted using the private key of the USS, which is assumed to stay private on the USS throughout the deployment of the system. If the private key of the USS is not leaked, based on the cryptographic properties of ECIES, a passive eavesdropper cannot obtain the plain-text location $\mathbf{x}$ used to generate the ciphertext $c$. The eavesdropper cannot even know if the plain-text location $\mathbf{x}$ changes or stays the same across consecutive location disclosures, thanks to the protection offered by ECIES against chosen-ciphertext, and chosen-plaintext attacks, as discussed in [ecies_standard].